OpenStudy (anonymous):

What is the equation of a line, in general form, that passes through points (-1, 2) and (5, 2)?

OpenStudy (shad0w):

This is the formula for finding the slope

OpenStudy (shad0w):

M is the slope

OpenStudy (anonymous):

& after that?

OpenStudy (anonymous):

\[\frac{y -y _{1} }{ x-x _{1} }=\frac{y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

\[\frac{y -2 }{ x+1}=\frac{2-2 }{ 5+1 }\] \[\frac{y -2 }{ x+1}=0\] \[y -2=0\] \[y =2\]

OpenStudy (anonymous):